# To Theoretically Understand Transformer-Based In-Context Learning for Optimizing CSMA

### v1

@@ 1. Introduction
There are also model-free studies proposing deep reinforcement learning (DRL) based approaches for unknown environments in the recent CSMA literature (e.g., (Wydmański and Szott, 2021; Yan et al., 2024; Lee et al., 2024)). (Wydmański and Szott, 2021) proposes a DRL approach to dynamically adjust contention window size based on turn-around-time measurement of channel status in IEEE 802.11ax networks. (Yan et al., 2024) presents a multi-device distributed DRL framework that intelligently tunes the contention window in IEEE 802.11bn networks to minimize tail latency while maintaining throughput. However, such approaches need to retrain from scratch once the channel environment changes, hard to implement on resource-constrained wireless devices. Though (Lee et al., 2024) proposes a soft actor-critic based approach for robust performance in response to environmental changes, it still inherits heavy training and inference costs. Thanks to the recent success of LLMs, a new idea naturally arises:
@@
In this paper, we study the theory of LLM transformer-based in-context learning (ICL) for optimizing channel access. ICL refers to the ability of a pretrained transformer to learn a new task simply by conditioning on a few input–output examples in its prompt, leveraging learned patterns and attention mechanisms without fine-tuning (e.g., (Dong et al., 2024), (Rubin et al., 2022), (Min et al., 2022b)). Unlike the costly DRL-based approaches, ICL operates through on-the-fly inference once pre-trained and is widely adaptable to many complex tasks like mathematical reasoning problems (e.g., (Liu et al., 2024), (Lu et al., [n. d.])). Recently, the communication society has explored the direction of introducing ICL to network design, such as transmission power allocation (e.g., (Lee and Park, 2024), (Zhou et al., 2024)), network deployment (e.g., (Boateng et al., 2024), (Sevim et al., 2024)), and network detection (e.g., (Song et al., 2024), (Zhang et al., 2024b)). However, all these studies are empirical, lacking a theoretical analysis with performance guarantees.
@@ 2. System Model and Problem Formulation
(1) $U=\frac{\mathbb{E}[\text{A successful packet transmission per slot}]}{\mathbb{E}[\text{Length of a slot time}]}.$
@@
For an improved design of ICL for channel access, a converged loss of our approach’s throughput $\hat{U}$ from that at the optimum $U^{*}$ should be guaranteed within a limited number of training steps. In the rest of this work, for a vector $w$, we let $\|w\|$ denote its $\ell$-222 norm. For some positive constant $c_{1}$ and $c_{2}$, we define $x=\Theta(y)$ if $c_{1}|y|<x<c_{2}|y|$ and $x=\mathcal{O}(y)$ if $x<c_{1}|y|$. Let $\mathbb{N}$ denote the set of natural numbers.
@@ 4.3. Step IV: Transformer Training
3:Record $\theta^{*}\leftarrow\theta^{(t+1)}$ if $\exists t<\bar{T}$ such that $||\theta^{(t+1)}-\theta^{(t)}||_{2}\leq\epsilon$. Otherwise, Record $\theta^{*}\leftarrow\theta^{(\bar{T})}$.
@@ 5.1. Convergence Analysis of Algorithm 1
$=$ $\frac{\exp\!\Bigl{(}\mathcal{M}\bigl{(}E^{x}(P^{s})\bigr{)}^{\!\top}\,Q\bigl{(}\theta^{(t)}\bigr{)}\Bigr{)}}{\displaystyle\sum_{j\in[M]}\exp\!\Bigl{(}\mathcal{M}\bigl{(}E^{x}_{j}(P^{s})\bigr{)}^{\top}\,Q\bigl{(}\theta^{(t)}\bigr{)}\Bigr{)}}.$
@@
Intuitively, in the convergence state, if the query token is $x_{q}^{s}=x_{k}$, the transformer correctly outputs the optimal contention window threshold $f^{s}(x_{k})$ for $x_{q}^{s}$ by assigning an attention score $\text{Attn}_{k}^{(t)}$ close to 111 in (11). Based on the convergence in attention score $\text{Attn}_{k}^{(t)}$ in Proposition 5.3, we now establish the convergence guarantee for Algorithm 1.
@@ 6.1. Extended System Model of the Erroneous Data Input
$\alpha\mathbb{P}_{f}(s_{2}|s_{1}\oplus``\mathrm{\backslash n}")\leq\mathbb{P}_{f}(s_{2})\leq\frac{1}{\alpha}\mathbb{P}_{f}(s_{2}|s_{1}\oplus``\mathrm{\backslash n}"),$ ” ) ,
@@ 6.2. Robustness Analysis to Erroneous Data Inputs
Suppose that the minimum KL-divergence between our ICL mapping $f$ and the ground-truth $f^{*}$ satisfies $\min_{f}KL(\mathbb{P}_{f},\mathbb{P}_{f^{*}})$$>-8\ln(\alpha\beta)$. If the number of in-context examples $M$ satisfies
@@ 7.2. Experimental Results in the Training Stage
[…]
@@ 7.4. More Practical Experiments on NS-3
For further comparison with recent CSMA advances, we now use NS-3.38 to resemble a real-world experiment. Wireless devices are randomly placed on a 20m-radius circle around a WiFi AP. The 802.11b link runs at 1 Mbps with 16 dBm transmit power and a 7 dB noise. Each device sends 1029-byte UDP packets every 1 ms to a server on the AP via port 8000. We use NS-3’s FlowMonitor to record the throughput as the total successfully received application-layer payload divided by the effective transmission time. We train our transformer only under small-scale node densities $N\in[2,6]$ with 20% erroneous prompts, optimize model-based approaches under an estimated node density of $N$=10, and optimize DRL-based approaches to adapt to unknown $N\in[50,150]$. We implement the Soft Actor–Critic (SAC) algorithm with both actor and twin-critic networks as two-layer MLPs (128 units per layer, ReLU activations), outputting a tanh-squashed Gaussian policy (log-std clamped to [20,2][-20,2][ - 20 , 2 ]). The critic and the actor are trained with Adam at a learning rate of $1\times 10^{-4}$, using a discount factor $\gamma=0.99$, soft-target updates ($\tau=0.005$), and automatic entropy tuning (initial $\alpha\approx 0.1$, targeting $-\text{dim}(\mathbf{a})$). We employ a small replay buffer (capacity 200020002000) to prioritize recent transitions, with full-batch updates every 20 steps followed by buffer clearance. Our ICL approach incurs the same learning rate.
Figure 6 shows prediction loss versus training steps for our ICL approach and soft actor-critic RL (the remaining DRL-based approaches are too slow to be comparable with ours), respectively. Figure 6 shows that the soft actor-critic RL incurs at least 650 training steps to achieve a stable convergence with $<$0.1 loss. In contrast, our approach attains zero loss within the first 100 steps, indicating a much faster convergence.
@@ 8. Conclusion
In this paper, we study LLM transformer-based ICL for optimizing channel access. We propose a transformer-based ICL optimizer to pre-collect collision-threshold data examples and a query collision case. They are constructed as a prompt as the input for the transformer to learn the pattern, which then generates a predicted CWT. To train the transformer for effective ICL, we develop an efficient algorithm and guarantee a near-optimal CWT prediction within limited training steps, ensuring a converged throughput loss. As it may be hard to gather perfect data examples in practice, we further extend to allow erroneous examples in the prompt. We prove that our optimizer still incurs limited prediction and throughput losses from the optimum. Experimental results on NS-3 further demonstrate our approach’s fast convergence and near-optimal throughput over existing model-based and DRL-based approaches under unknown node densities.
@@
To prove the lower bound in Theorem 3.1, we first prove that there exists a $C_{1}>0$ such that $|\tau(N)-\tau(\hat{N})|\geq C_{1}|N-\hat{N}|$. Then, we prove that there exists a $C_{2}>0$ such that $|U(\tau(N))-U(\tau(\hat{N}))|\geq C_{2}|\tau(N)-\tau(\hat{N})|$.
@@
$|\tau(N)-\tau(\hat{N})|=\bigg{|}\frac{\partial\tau}{\partial N}\big{|}_{\tau=\tau_{0}}\bigg{|}|N-\hat{N}|.$
Therefore, to prove that $|\tau(N)-\tau(\hat{N})|\geq C_{1}|N-\hat{N}|$, it is enough to prove that $|\frac{\partial\tau}{\partial N}|\geq C_{1}$. This motivates us to find a lower bound of the numerator of $|\frac{\partial\tau}{\partial N}|$ and an upper bound of the denominator of $|\frac{\partial\tau}{\partial N}|$. The partial derivative of $D(N,\tau)$ on $N$ is:
@@
$|U(\tau(n))-U(\tau(\hat{n}))|\geq\frac{T_{\sigma}}{T_{c}^{2}}|\tau(n)-\tau(\hat{n})|\geq\Theta\bigg{(}\frac{T_{\sigma}}{\bar{N}K^{2}\bar{W}^{3}T_{c}^{2}}\bigg{)}|N-\hat{N}|.$
@@
(23) :=$\mathcal{E}^{*}:=\left\{P^{(t)}_{1:N}:|\mathcal{X}^{s}_{k}|\in\left[p_{k}N-\frac{cN}{K},p_{k}N+\frac{cN}{K}\right]\text{ for }k\in[K]\right\}.$
@@
For ease of exposition, we define $u_{k}:=(p_{k}-\delta_{0})K$ and $U_{k}:=(p_{k}+\delta_{0})K$, which satisfy $u_{k}=\Theta(1)$ and $U_{k}=\Theta(1)$, given $p_{k}=\Theta(\frac{1}{K})$ and $\delta_{0}=\mathcal{O}(\frac{1}{K})$. Then for any $P_{1:N}^{(t)}$ belonging to $\mathcal{E}^{*}$, we have $|\mathcal{X}_{k}|\in[\frac{u_{k}N}{K},\frac{UK}{N}]=\Theta(\frac{N}{K})$ with a probability close to 111. We then prove Proposition 5.3 based on Lemma C.1.
@@ D.2. Convergence of prediction loss $\mathcal{L}(\theta)$
[…]
@@
Suppose that the minimum KL-divergence between our ICL mapping $f$ and the ground-truth $f^{*}$ satisfies $\min_{f}KL(\mathbb{P}_{f},\mathbb{P}_{f^{*}})$$>-8\ln(\alpha\beta)$. If the number of in-context data examples $M$ is long enough as
@@
(33) $\mathbb{P}_{f}(x_{m}\oplus W_{m})\geq\mathbb{P}_{f}(x_{m}\oplus W_{m}\oplus``\mathrm{\backslash n}")=\mathbb{P}_{f}(x_{m}\oplus W_{m})\cdot\mathbb{P}_{f}(``\mathrm{\backslash n}"|x_{m}\oplus W_{m})>\mathbb{P}_{f}(x_{m}\oplus W_{m})\cdot\beta,$
@@
(35) $\beta<\frac{\mathbb{P}_{f}(x_{m}\oplus W_{m})}{\mathbb{P}_{f}(x_{m}\oplus W_{m}\oplus``\mathrm{\backslash n}")}<\beta^{-1}.$
@@
[…]
@@
[…]
@@
[…]
@@
$\mathbb{E}[U(W)-U(\hat{W})]=$ $\mathbb{E}[|U(W)-U(\hat{W}|]$
$\leq$ $\mathbb{E}\bigg{[}\frac{T_{P}\bar{N}}{8T_{\sigma}}\cdot|W-\hat{W}|\bigg{]}$

### v4

@@ Abstract.
This work is supported in part by SUTD Kickstarter Initiative (SKI) Grant with no. SKI 2021_04_07; and in part by the Joint SMU-SUTD Grant with no. 22-LKCSB-SMU-053.
@@ 1. Introduction
There are also model-free studies proposing deep reinforcement learning (DRL) based approaches for unknown environments in the recent CSMA literature (e.g., (Wydmański and Szott, 2021; Yan et al., 2024; Lee et al., 2024)). (Wydmański and Szott, 2021) proposes a DRL approach to dynamically adjust contention window size based on turn-around-time measurement of channel status in IEEE 802.11ax networks. (Yan et al., 2024) presents a multi-device distributed DRL framework that intelligently tunes the contention window in IEEE 802.11bn networks to minimize tail latency while maintaining throughput. However, such approaches need to retrain from scratch once the channel environment changes, hard to implement on resource-constrained wireless devices. Though (Lee et al., 2024) proposes a soft actor-critic based approach for robust performance in response to environmental changes, it still inherits heavy training and inference costs. Thanks to the recent success of AI/LLMs (e.g., (Li and Duan, 2025),(Li et al., [n. d.])), a new idea arises:
@@
In this paper, we study the theory of LLM transformer-based in-context learning (ICL) for optimizing channel access. ICL refers to the ability of a pretrained transformer to learn a new task simply by conditioning on a few input–output examples in its prompt, leveraging learned patterns and attention mechanisms without fine-tuning (e.g., (Dong et al., 2024), (Rubin et al., 2022), (Min et al., 2022b)). Unlike the costly DRL-based approaches, ICL operates through on-the-fly inference once pre-trained and is widely adaptable to many complex tasks like mathematical reasoning problems (e.g., (Liu et al., 2024), (Lu et al., [n. d.])). Recently, the communication society has explored the direction of introducing ICL to network design, such as transmission power allocation (e.g., (Lee and Park, 2024), (Zhou et al., 2024)), network deployment (e.g., (Boateng et al., 2024), (Sevim et al., 2024),(Hao and Duan, 2025b),(Hao and Duan, 2025a)), and network detection (e.g., (Song et al., 2024), (Zhang et al., 2024b)). However, all these studies are empirical, lacking a theoretical analysis with performance guarantees.
@@ 2. System Model and Problem Formulation
[…]
@@
For an improved design of ICL for channel access, a converged loss of our approach’s throughput $\hat{U}$ from that at the optimum $U^{*}$ should be guaranteed within a limited number of training steps. In the rest of this work, for a vector $w$, we let $\|w\|$ denote its $\ell$-22 norm. For some positive constant $c_{1}$ and $c_{2}$, we define $x=\Theta(y)$ if $c_{1}|y|<x<c_{2}|y|$ and $x=\mathcal{O}(y)$ if $x<c_{1}|y|$. Let $\mathbb{N}$ denote the set of natural numbers.
@@ 4.3. Step IV: Transformer Training
3:Record $\theta^{*}\leftarrow\theta^{(t+1)}$ if $\exists t<\bar{T}$ such that $\theta^{(t+1)}-\theta^{(t)}||_{2}\leq\epsilon$. Otherwise, Record $\theta^{*}\leftarrow\theta^{(\bar{T})}$.
@@ 5.1. Convergence Analysis of Algorithm 1
$=$ $\frac{\exp\!\Bigl{(}\bigl{(}E^{x}_{m}(P^{s})\bigr{)}^{\!\top}\,Q\bigl{(}\theta^{(t)}\bigr{)}E^{x}_{M+1}(P^{s})\Bigr{)}}{\displaystyle\sum_{j\in[M]}\exp\!\Bigl{(}\bigl{(}E^{x}_{j}(P^{s})\bigr{)}^{\top}\,Q\bigl{(}\theta^{(t)}\bigr{)}E^{x}_{M+1}(P^{s})\Bigr{)}}.$
@@
Intuitively, in the convergence state, if the query token is $x_{q}^{s}=x_{k}$, the transformer correctly outputs the optimal contention window threshold $f^{s}(x_{k})$ for $x_{q}^{s}$ by assigning an attention score $\text{Attn}_{k}^{(t)}$ close to $1$ in (11). Based on the convergence in attention score $\text{Attn}_{k}^{(t)}$ in Proposition 5.3, we now establish the convergence guarantee for Algorithm 1.
@@ 6.1. Extended System Model of the Erroneous Data Input
[…]
@@ 6.2. Robustness Analysis to Erroneous Data Inputs
Suppose that the minimum KL-divergence between our ICL mapping $f$ and the ground-truth $f^{*}$ satisfies $\min_{f}KL(\mathbb{P}_{f},\mathbb{P}_{f^{*}})$$8\ln(\alpha\beta)$. If the number of in-context examples $M$ satisfies
@@ 7.2. Experimental Results in the Training Stage
We consider node density environments $N\in$$\{2,3,4,5,6\}$ for ICL data collection. As in Step I of Figure 2, for each node density $N$, the AP system offline collects $M=9$ data points $\{(x_{k},W_{k}^{*})\}_{k=0}^{8}$ to construct the prompt, where $x_{k}=(k,T_{P},T_{s},T_{c})$, $k\in[K]$ denotes the current collision number and the maximum collision number is set as $K=8$. Note that the node density $N$ is not shown in any $x_{k}$. To obtain the contention window thresholds $\{W_{k}^{*}\}_{k=0}^{8}$ in a prompt with a node density $N$, the system first maximizes the throughput $U$ in (3) for obtaining the optimal packet transmission probability $\tau^{*}$ (e.g., using the branch-and-bound algorithm). Then, it uses the $\tau^{*}$ to solve $\{W_{k}^{*}\}_{k=0}^{8}$ according to (2) by applying MILP algorithms like golden section search.
As in Step II of Figure 2, we construct $S=5$ prompts corresponding to 5 different node densities $N\in$$\{2,3,4,5,6\}$ for training the transformer parameters $\theta$, where each prompt contains 8 example points $(x_{k},W_{k}^{*})$ and a query collision case $x_{q}^{s}$. These prompts are then embedded as in Step III of Figure 2. Regarding Step IV of Figure 2 for transformer training, we set the step size of Algorithm 1 as $\eta=0.05$. We consider $x_{q}^{s}=x_{k}$ to be the same for all 5 prompts and obtain our ICL prediction $\hat{W}_{k}=\hat{W}_{q}^{s}$ in (8) of each contention window threshold $W_{k}^{*}$, $k\in[K]$. Then, wireless devices download the collision-threshold pair table to use online for CSMA, fully compatible with legacy devices.
@@ 7.4. More Practical Experiments on NS-3
For further comparison with recent CSMA advances, we now use NS-3.38 to resemble a real-world experiment. Wireless devices are randomly placed on a 20m-radius circle around a WiFi AP. The 802.11b link runs at 1 Mbps with 16 dBm transmit power and a 7 dB noise. Each device sends 1029-byte UDP packets every 1 ms to a server on the AP via port 8000. We use NS-3’s FlowMonitor to record the throughput as the total successfully received application-layer payload divided by the effective transmission time. We train our transformer only under small-scale node densities $N\in[2,6]$ with 20% erroneous prompts, optimize model-based approaches under an estimated node density of NN=10, and optimize DRL-based approaches to adapt to unknown $N\in[50,150]$. We implement the Soft Actor–Critic (SAC) algorithm with both actor and twin-critic networks as two-layer MLPs (128 units per layer, ReLU activations), outputting a tanh-squashed Gaussian policy (log-std clamped to [20,2][-20,2]). The critic and the actor are trained with Adam at a learning rate of $1\times 10^{-4}$, using a discount factor $\gamma=0.99$, soft-target updates ($\tau=0.005$), and automatic entropy tuning (initial $\alpha\approx 0.1$, targeting $\text{dim}(\mathbf{a})$). We employ a small replay buffer (capacity $2000$) to prioritize recent transitions, with full-batch updates every 20 steps followed by buffer clearance. Our ICL approach incurs the same learning rate.
Figure 6 shows prediction loss versus training steps for our ICL approach and soft actor-critic RL (the remaining DRL-based approaches are too slow to be comparable with ours), respectively. Figure 6 shows that the soft actor-critic RL incurs at least 650 training steps to achieve a stable convergence with <<0.1 loss. In contrast, our approach attains zero loss within the first 100 steps, indicating a much faster convergence.
@@ 8. Conclusion
In this paper, we study LLM transformer-based ICL for optimizing channel access. We propose a transformer-based ICL optimizer to pre-collect collision-threshold data examples and a query collision case. They are constructed as a prompt as the input for the transformer to learn the pattern, which then generates a predicted CWT. To train the transformer for effective ICL, we develop an efficient algorithm and guarantee a near-optimal CWT prediction within limited training steps, ensuring a converged throughput loss. As it may be hard to gather perfect data examples in practice, we further extend to allow erroneous examples in the prompt. We prove that our optimizer still incurs limited prediction and throughput losses from the optimum. Experimental results on NS-3 further demonstrate our approach’s fast convergence and near-optimal throughput over existing model-based and DRL-based approaches.
@@
[…]
@@
$N^|.\displaystyle|\tau(N)-\tau(\hat{N})|=\bigg{|}\frac{\partial\tau}{\partial N}\big{|}_{\tau=\tau_{0}}\bigg{|}$|N-$\hat${N}|.
Therefore, to prove that $N^||\tau(N)-\tau(\hat{N})|\geq C_{1}$|N-$\hat${N}|, it is enough to prove that $|\frac{\partial\tau}{\partial N}|\geq C_{1}$. This motivates us to find a lower bound of the numerator of $|\frac{\partial\tau}{\partial N}|$ and an upper bound of the denominator of $|\frac{\partial\tau}{\partial N}|$. The partial derivative of $D(N,\tau)$ on $N$ is:
@@
$N^|.\displaystyle|U(\tau(n))-U(\tau(\hat{n}))|\geq\frac{T_{\sigma}}{T_{c}^{2}}|\tau(n)-\tau(\hat{n})|\geq\Theta\bigg{(}\frac{T_{\sigma}}{\bar{N}K^{2}\bar{W}^{3}T_{c}^{2}}\bigg{)}$|N-$\hat${N}|.
@@
(23) $\mathcal{E}^{*}:=\left\{P^{(t)}_{1:N}:|\mathcal{X}^{s}_{k}|\in\left[p_{k}N-\frac{cN}{K},p_{k}N+\frac{cN}{K}\right]\text{ for }k\in[K]\right\}.$
@@
For ease of exposition, we define $u_{k}:=(p_{k}-\delta_{0})K$ and $U_{k}:=(p_{k}+\delta_{0})K$, which satisfy $u_{k}=\Theta(1)$ and $U_{k}=\Theta(1)$, given $p_{k}=\Theta(\frac{1}{K})$ and $\delta_{0}=\mathcal{O}(\frac{1}{K})$. Then for any $P_{1:N}^{(t)}$ belonging to $\mathcal{E}^{*}$, we have $|\mathcal{X}_{k}|\in[\frac{u_{k}N}{K},\frac{UK}{N}]=\Theta(\frac{N}{K})$ with a probability close to $1$. We then prove Proposition 5.3 based on Lemma C.1.
@@ D.2. Convergence of prediction loss $\mathcal{L}(\theta)$
where the first inequality holds due to $(1-(1-\tau)^{N-1})^{k}\leq 1$, the second due to $\tau(1-\tau)^{N-1}$ increasing with $\tau\leq\frac{1}{N}$ (this bound $1/N$ comes from Lemma A.1), and the last due to $\frac{1}{N}(1-\frac{1}{N})^{N-1}$ decreasing with $N\geq 2$. Further, we have
@@
Suppose that the minimum KL-divergence between our ICL mapping $f$ and the ground-truth $f^{*}$ satisfies $\min_{f}KL(\mathbb{P}_{f},\mathbb{P}_{f^{*}})$$8\ln(\alpha\beta)$. If the number of in-context data examples $M$ is long enough as
@@
[…]
@@
[…]
@@
Denote $P=x_{1}\oplus W_{1}\oplus``\mathrm{\backslash n}"\oplus\cdots\oplus x_{M}\oplus W_{M}\oplus``\mathrm{\backslash n}$". We have
@@
Denote $P^{\prime}=x_{1}\oplus W_{1}\oplus``\mathrm{\backslash n}"\oplus\cdots\oplus x_{M}\oplus W_{M}\oplus``\mathrm{\backslash n}$" and $P=P^{\prime}\oplus x$. We have
@@
due to $Pr(f|\mathcal{D^{*}})>\beta^{\ell}$ and $Pr(f|\mathcal{D^{*}})\geq\gamma$. According to Lemma G.1, as long as $\frac{\mathbb{P}_{f}(P^{\prime})}{\mathbb{P}_{f^{*}}(P^{\prime})}\leq\frac{\mathbb{P}_{\mathcal{D^{*}}}(W|x)-\mathbb{P}_{\mathcal{D^{*}}}(\hat{W}|x)}{5\alpha^{-2}\beta^{-\ell}\gamma^{-1}}:=\mu$ and the $M$ is properly choosen, we have $\bigg{|}\frac{B}{C}\bigg{|}\leq\frac{1}{5}(\mathbb{P}_{\mathcal{D^{*}}}(W|x)-\mathbb{P}_{\mathcal{D^{*}}}(\hat{W}|x))$ with probability at least 1q1-q. Similarly, we can bound $|\frac{D}{C}|<\frac{1}{4}$.
@@
Next, we prove our throughput loss bound. We have the ICL prediction loss $Pr(\mathbb{E}[\mathbf{1}_{(\hat{W}\neq W)}]\leq\epsilon+\text{BER})\geq 1-q$, where BER stands for error rate of the Bayes optimal classifier. With probability of at least 1q1-q, We have
@@
[…]